# Ophiuchus: Scalable Modeling of Protein Structures through Hierarchical Coarse-graining SO(3)-Equivariant Autoencoders

### v1

@@ Abstract
Three-dimensional native states of natural proteins display recurring and hierarchical patterns. Yet, traditional graph-based modeling of protein structures is often limited to operate within a single fine-grained resolution, and lacks hourglass neural architectures to learn those high-level building blocks. We narrow this gap by introducing Ophiuchus, an SO(3)-equivariant coarse-graining model that efficiently operates on all heavy atoms of standard protein residues, while respecting their relevant symmetries. Our model departs from current approaches that employ graph modeling, instead focusing on local convolutional coarsening to model sequence-motif interactions in log-linear length complexity. We train Ophiuchus on contiguous fragments of PDB monomers, investigating its reconstruction capabilities across different compression rates. We examine the learned latent space and demonstrate its prompt usage in conformational interpolation, comparing interpolated trajectories to structure snapshots from the PDBFlex dataset. Finally, we leverage denoising diffusion probabilistic models (DDPM) to efficiently sample readily-decodable latent embeddings of diverse miniproteins. Our experiments demonstrate Ophiuchus to be a scalable basis for efficient protein modeling and generation.
@@ 1 Introduction
Proteins form the basis of all biological processes and understanding them is critical to biological discovery, medical research and drug development. Fortunately, their three-dimensional structures often display modular organization across multiple scales, making them promising candidates for modeling in motif-based design spaces [Bystroff & Baker (1998); Mackenzie & Grigoryan (2017); Swanson et al. (2022)]. Harnessing these coarser, lower-frequency building blocks is of great relevance to the investigation of the mechanisms behind protein evolution, folding and dynamics [Mackenzie et al. (2016)], and may be instrumental in enabling more efficient computation on protein structural data through coarse and latent variable modeling [Kmiecik et al. (2016); Ramaswamy et al. (2021)].
Deep learning models have been shown to produce representations that capture high-level patterns of the data [Olah et al. (2018)]. Recent developments in learning architectures to protein structures demonstrates the remarkable capabilities of neural models in the domain of protein modeling and design [Jumper et al. (2021); Baek et al. (2021b); Ingraham et al. (2022); Watson et al. (2022)]. Still, current state-of-the-art learning architectures often fail to incorporate inductive biases necessary for directly representing high-level components, such as coarsening of representations over the data domain.
To fill this gap, we introduce Ophiuchus, a deep convolutional model that captures joint encodings of sequence-structure motifs of all-atom protein structures. Our model contracts representations over the sequence to produce compressed encodings in hierarchical fashion. To capture three-dimensional geometry effectively, we model features as irreducible representations of SO(3), introducing novel roto-translational equivariant neural architectures for multi-layer hourglass-shaped autoencoder.
Figure 1: Coarsening a Three-Dimensional Sequence: (a) Ophiuchus takes atoms directly as representations of SO(3). Non-scalar signals are taken relative to the $\mathbf{C}_{\alpha}$ position of the corresponding residue. Only 3D vector $\mathbf{V}^{l=1}$ components are shown. ($\ast$) Unorderable atoms are encoded in a permutation invariant way. (b) Our proposed model uses three building blocks to transform those vectors and their anchors into coarser representations of coordinates. (c) We train a deep bottlenecking model by iteratevily coarsening and refining across the sequence axis.
We ablate the reconstruction performance of different network layouts and compression settings. We investigate the learned latent space and show it to produce rich representations that are directly employable in structural interpolation. Finally, we deploy a diffusion model on the latent space of Ophiuchus to investigate the efficacy of a coarse-to-fine method to protein structures.
Our main contributions are the following:
  • Deep Protein Autoencoding: We introduce a new autoencoding model for protein sequence-structure representation, and offer an in-depth examination of its architecture through comprehensive ablation benchmarks. We showcase the unique effectiveness and versatility of our model’s latent geometric representations in structural latent interpolation tasks.
  • Neural Architecture for Proteins: We propose novel learning algorithms for efficiently operating on all standard protein heavy atoms while preserving the natural permutation invariance of particular side-chain atoms. We introduce numerous methods for coarsening, refining and mixing of protein sequence-structure, employing irreducible representations of SO(3) to contract and spawn 3D coordinates directly from features.
@@
  • All-Atom Protein Latent Diffusion: We explore a novel generative approach to proteins by performing latent diffusion exclusively on geometric feature representations. We train an efficient all-atom generative model for miniproteins, and provide diverse benchmarks to assess its performance.
@@ 2.1 Modularity and Hierarchy in Proteins
Protein sequences and structures display significant degrees of modularity. [Vallat et al. (2015)] introduces a library of common super-secondary structural motifs (Smotifs), while [Mackenzie et al. (2016)] shows protein structural space to be efficiently describable by small tertiary alphabets (TERMs). Motif-based methods have been successfully used in protein folding and design [Bystroff & Baker (1998); Li et al. (2022)].
@@ 2.2 Symmetries in Neural Architecture for Biomolecules
Translation Equivariance of Sequence: One-dimensional Convolutional Neural Networks (CNNs) have been demonstrated to successfully model protein sequences across a variety of tasks Karydis (2017); Hou et al. (2018); Lee et al. (2019); Yang et al. (2022). These models capture sequence-motif representations that are equivariant to translation of the sequence. However, sequential convolution is less common in architectures for protein structures, which are often cast as Graph Neural Networks (GNNs) [Zhang et al. (2021)]. Notably, [Fan et al. (2022)] proposes a CNN network to model the regularity of one-dimensional sequences along with three-dimensional structures, but they restrict their layout to coarsening. In this work, we further integrate geometry into sequence by directly using three-dimensional vector feature representations and transformations in 1D convolutions. We use this CNN to investigate an unsupervised autoencoding approach to protein structures.
Permutation Invariances of Side-chain Atoms. In order to capture the permutable ordering of atoms, neural models of molecules are often implemented with permutation-invariant GNNs [Wieder et al. (2020)]. Nevertheless, protein structures are sequentially ordered, and most standard side-chain atoms are readily orderable, with four exceptions of residues [Jumper et al. (2021)]. We use this fact to design an efficient approach to directly model all heavy atoms of proteins, proposing a method to parse atomic positions and symmetries through roto-translational equivariant features.
@@ 2.3 Unsupervised Learning of Proteins
Unsupervised techniques for capturing protein sequence and structure have witnessed remarkable advancements in recent years [Lin et al. (2023); Elnaggar et al. (2023); Zhang et al. (2022)]. Amongst unsupervised methods, autoencoder models learn to produce efficient low-dimensional representations using an informational bottleneck. These models have received significant attention for enabling learning of coarse representations of molecules [Wang & Gómez-Bombarelli (2019); Yang & Gómez-Bombarelli (2023); Winter et al. (2021); Wehmeyer & Noé (2018); Ramaswamy et al. (2021)], and have been successfully deployed to diverse protein tasks of modeling and sampling [Wang et al. (2022); Eguchi et al. (2020); Lin et al. (2021); Mansoor et al. (2023)].
@@ 2.4 Denoising Diffusion for Proteins
Denoising Diffusion Probabilistic Models (DDPMs) [Sohl-Dickstein et al. (2015); Ho et al. (2020)] have found widespread adoption through diverse architectures for generative sampling of protein structures. Chroma [Ingraham et al. (2022)] trains random graph message passing through roto-translational invariant features, while RFDiffusion [Watson et al. (2022)] fine-tunes pretrained folding model RoseTTAFold [Baek et al. (2021a)] to denoising, employing SE(3)-equivariant transformers in structural decoding [Fuchs et al. (2020)]. [Yim et al. (2023); Anand & Achim (2022)] generalize denoising diffusion to frames of reference, employing Invariant Point Attention [Jumper et al. (2021)] to model three dimensions. More recently, [Fu et al. (2023)] designs a latent diffusion model by employing EGNNs [Satorras et al. (2022)] to contract the coordinate dimension into a smaller space. In contrast, our model uses SO(3)-equivariant feature representations to produce increasingly richer structural features from autoencoded coordinates. We perform generation directly on these features to demonstrate a latent diffusion generative model.
@@ 3 The Ophiuchus Architecture
We treat a protein as an ordered list $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})_{1:N}$ of $N$ nodes, each with an anchoring position $\mathbf{P}_{i}\in\mathbb{R}^{1\times 3}$ and a collection of irreducible representations of SO(3): $\mathbf{V}_{i}^{l}\in\mathbb{R}^{d\times(2l+1)}$ for integers $l\in[0,{l_{\max}}]$, $i\in[1,N]$. These representations are directly produced from sequence labels and all-atom positions, as we describe in the Atom Encoder/Decoder sections. Our model then acts with diverse operations on these representations. Self-Interaction updates vector representations independently, mixing within same-anchored vectors of different orders. In order to mix across the domain, Sequence Convolution updates $K$-sized windows of the list simultaneously, message-passing within sequence neighbors. Finally, Spatial Convolutions employ permutation invariant geometric messages to update the representations and coordinates of spatial neighbors. These modules are composed to build a coarsening-refining encoder-decoder hourglass architecture.
@@ 3.1 All-Atom Atom Encoder and Decoder
The input format of our modeling starts at the residue level. Given a particular $i$-th residue, let $\mathbf{R}_{i}\in\mathcal{R}$ denote its residue label, $\mathbf{P}_{i}^{\alpha}\in\mathbb{R}^{1\times 3}$ denote the global position of its alpha carbon ($\mathbf{C}_{\alpha}$), and $\mathbf{P}^{\ast}_{i}$ the position of all other heavy atoms relative to $\mathbf{C}_{\alpha}$. We produce initial residue representations $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})^{0}_{i}$ by setting anchors $\mathbf{P}_{i}^{0}=\mathbf{P}^{\alpha}_{i}$, scalar representations $\mathbf{P}_{i}^{0,l=0}=\textrm{Embed}(\mathbf{R}_{i})$ , and vector representations $\mathbf{V}_{i}^{0,l>0}$ to explicitly encode relative atomic positions $\mathbf{P}^{\ast}_{i}$. In particular, provided the residue label $\mathbf{R}_{i}$, heavy atoms of most standard protein residues are readily orderable and may be directly processed as a stack of $\mathbf{V}^{l=1}$ relative vector positions.
However, some of the standard residues do present two-permutations within their atom configurations (i.e. they are not orderable). To handle these cases, we map the set of two relative vectors into a pair $(\mathbf{V}^{l=1},\mathbf{V}^{l=2})$ where $\mathbf{V}^{l=1}$ is the center and $\mathbf{V}^{l=2}$ is the unsigned difference. This signal is invariant to corresponding atomic two-flips, while still carrying relevant information on positioning and angularity. Decoding is handled similarly. Refer to Appendix A for further details.
Figure 2: Building Blocks of Ophiuchus: (a) Atom Encoder and (b) Atom Decoder enable the model to directly take and produce atomic coordinates. (c) Self Interaction updates representations internally, across different vector orders $l$. (d) Sequence Convolution and (e) Transpose Sequence Convolution communicate sequence neighbors and produce coarser and finer representations, respectively. (d) Spatial Convolution interacts spatial neighbors. (g) Hourglass model: we compose those modules to build encoder and decoder models, stacking them into an autoencoder.
@@
Input: Heavy Atom Relative Positions $\mathbf{P}^{\ast}\in\mathbb{R}^{n\times 3}$
@@
$\mathbf{P}\leftarrow\mathbf{P}^{\alpha}$ $\mathbf{V}^{l=0}\leftarrow\textrm{Embed}(\mathbf{R})$ $\mathbf{V}^{l=1}\leftarrow\textrm{UnorderableCenters}(\mathbf{R},\mathbf{P}^{\ast})$ $\mathbf{V}^{l=2}\leftarrow\textrm{UnorderableDifferences}(\mathbf{R},\mathbf{P}^{\ast})$ $\mathbf{V}^{l=1}\leftarrow\mathbf{V}^{l=1}\oplus\textrm{Orderable}(\mathbf{R},\mathbf{P}^{\ast})$ return $(\mathbf{P},\mathbf{V}^{l=0:2})$
@@
Output: Heavy Atom Relative Positions $\mathbf{P}^{\ast}\in\mathbb{R}^{n\times 3}$
Output: Residue Label $\mathbf{R}\in\mathcal{R}$
$\mathbf{P}^{\alpha}\leftarrow\mathbf{P}$ $\mathbf{R}\leftarrow\mathrm{Softmax}(\mathbf{V}^{l=0})$ $\mathbf{P}^{\ast}\leftarrow\mathrm{Orderable}(\mathbf{R},\mathbf{V}^{l=1})$ $\mathbf{P}^{\ast}\leftarrow\mathbf{P}^{\ast}\oplus\mathrm{Unorderable}(\mathbf{R},\mathbf{V}^{l=1},\mathbf{V}^{l=2})$ return $(\mathbf{P}^{\alpha},\mathbf{P}^{\ast},\mathbf{R})$
@@
This processing makes Ophiuchus strictly blind to ordering flips within permutable atoms, while still enabling it to operate directly on all heavy atoms in an efficient, stacked representation.

### v2

@@ Abstract
Three-dimensional native states of natural proteins display recurring and hierarchical patterns. Yet, traditional graph-based modeling of protein structures is often limited to operate within a single fine-grained resolution, and lacks hourglass neural architectures to learn those high-level building blocks. We narrow this gap by introducing Ophiuchus, an SO(3)-equivariant coarse-graining model that efficiently operates on all-atom protein structures. Our model departs from current approaches that employ graph modeling, instead focusing on local convolutional coarsening to model sequence-motif interactions with efficient time complexity in protein length. We measure the reconstruction capabilities of Ophiuchus across different compression rates, and compare it to existing models. We examine the learned latent space and demonstrate its utility through conformational interpolation. Finally, we leverage denoising diffusion probabilistic models (DDPM) in the latent space to efficiently sample protein structures. Our experiments demonstrate Ophiuchus to be a scalable basis for efficient protein modeling and generation.
@@ 1 Introduction
Proteins form the basis of all biological processes and understanding them is critical to biological discovery, medical research and drug development. Their three-dimensional structures often display modular organization across multiple scales, making them promising candidates for modeling in motif-based design spaces [Bystroff & Baker (1998); Mackenzie & Grigoryan (2017); Swanson et al. (2022)]. Harnessing these coarser, lower-frequency building blocks is of great relevance to the investigation of the mechanisms behind protein evolution, folding and dynamics [Mackenzie et al. (2016)], and may be instrumental in enabling more efficient computation on protein structural data through coarse and latent variable modeling [Kmiecik et al. (2016); Ramaswamy et al. (2021)].
Recent developments in deep learning architectures applied to protein sequences and structures demonstrate the remarkable capabilities of neural models in the domain of protein modeling and design [Jumper et al. (2021); Baek et al. (2021b); Ingraham et al. (2022); Watson et al. (2022)]. Still, current state-of-the-art architectures lack the structure and mechanisms to directly learn and operate on modular protein blocks.
To fill this gap, we introduce Ophiuchus, a deep SO(3)-equivariant model that captures joint encodings of sequence-structure motifs of all-atom protein structures. Our model is a novel autoencoder that uses one-dimensional sequence convolutions on geometric features to learn coarsened representations of proteins. Ophiuchus outperforms existing SO(3)-equivariant autoencoders [Fu et al. (2023)] on the protein reconstruction task. We present extensive ablations of model performance across different autoencoder layouts and compression settings. We demonstrate that our model learns a robust and structured representation of protein structures by learning a denoising diffusion probabilistic model (DDPM) [Ho et al. (2020)] in the latent space. We find Ophiuchus to enable significantly faster sampling of protein structures, as compared to existing diffusion models [Wu et al. (2022a); Yim et al. (2023); Watson et al. (2023)], while producing unconditional samples of comparable quality and diversity.
Figure 1: Coarsening a Three-Dimensional Sequence. (a) Each residue is represented with its $\mathbf{C}_{\alpha}$ atom position and geometric features that encode its label and the positions of its other atoms. ($\ast$) side-chains with non-orderable atoms are encoded in a permutation-invariant way. (b) Our proposed model uses roto-translation equivariant convolutions to coarsen these positions and geometric features. (c) We train deep autoencoders to reconstruct all-atom protein structures directly in three-dimensions.
Our main contributions are summarized as follows:
  • Novel Autoencoder: We introduce a novel SO(3)-equivariant autoencoder for protein sequence and all-atom structure representation. We propose novel learning algorithms for coarsening and refining protein representations, leveraging irreducible representations of SO(3) to efficiently model geometric information. We demonstrate the power of our latent space through unsupervised clustering and latent interpolation.
  • Extensive Ablation: We offer an in-depth examination of our architecture through extensive ablation across different protein lengths, coarsening resolutions and model sizes. We study the trade-off of producing a coarsened representation of a protein at different resolutions and the recoverability of its sequence and structure.
@@
  • Latent Diffusion: We explore a novel generative approach to proteins by performing latent diffusion on geometric feature representations. We train diffusion models for multiple resolutions, and provide diverse benchmarks to assess sample quality. To the best of our knowledge, this is the first generative model to directly produce all-atom structures of proteins.
@@ 2.1 Modularity and Hierarchy in Proteins
Protein sequences and structures display significant degrees of modularity. [Vallat et al. (2015)] introduces a library of common super-secondary structural motifs (Smotifs), while [Mackenzie et al. (2016)] shows protein structural space to be efficiently describable by small tertiary alphabets (TERMs). Motif-based methods have been successfully used in protein folding and design [Bystroff & Baker (1998); Li et al. (2022)]. Inspired by this hierarchical nature of proteins, our proposed model learns coarse-grained representations of protein structures.
@@ 2.2 Symmetries in Neural Architecture for Biomolecules
Translation Equivariance of Sequence. One-dimensional Convolutional Neural Networks (CNNs) have been demonstrated to successfully model protein sequences across a variety of tasks [Karydis (2017); Hou et al. (2018); Lee et al. (2019); Yang et al. (2022)]. These models capture sequence-motif representations that are equivariant to translation of the sequence. However, sequential convolution is less common in architectures for protein structures, which are often cast as Graph Neural Networks (GNNs) [Zhang et al. (2021)]. Notably, [Fan et al. (2022)] proposes a CNN network to model the regularity of one-dimensional sequences along with three-dimensional structures, but they restrict their layout to coarsening. In this work, we further integrate geometry into sequence by directly using three-dimensional vector feature representations and transformations in 1D convolutions. We use this CNN to investigate an autoencoding approach to protein structures.
Permutation Invariances of Atomic Order. In order to capture the permutable ordering of atoms, neural models of molecules are often implemented with permutation-invariant GNNs [Wieder et al. (2020)]. Nevertheless, protein structures are sequentially ordered, and most standard side-chain heavy atoms are readily orderable, with exception of four residues [Jumper et al. (2021)]. We use this fact to design an efficient approach to directly model all-atom protein structures, introducing a method to parse atomic positions in parallel channels as roto-translational equivariant feature representations.
@@ 2.3 Unsupervised Learning of Proteins
Unsupervised techniques for capturing protein sequence and structure have witnessed remarkable advancements in recent years [Lin et al. (2023); Elnaggar et al. (2023); Zhang et al. (2022)]. Amongst unsupervised methods, autoencoder models learn to produce efficient low-dimensional representations using an informational bottleneck. These models have been successfully deployed to diverse protein tasks of modeling and sampling [Eguchi et al. (2020); Lin et al. (2021); Wang et al. (2022); Mansoor et al. (2023); Visani et al. (2023)], and have received renewed attention for enabling the learning of coarse representations of molecules [Wang & Gómez-Bombarelli (2019); Yang & Gómez-Bombarelli (2023); Winter et al. (2021); Wehmeyer & Noé (2018); Ramaswamy et al. (2021)]. However, existing three-dimensional autoencoders for proteins do not have the structure or mechanisms to explore the extent to which coarsening is possible in proteins. In this work, we fill this gap with extensive experiments on an autoencoder for deep protein representation coarsening.
@@ 2.4 Denoising Diffusion for Proteins
Denoising Diffusion Probabilistic Models (DDPM) [Sohl-Dickstein et al. (2015); Ho et al. (2020)] have found widespread adoption through diverse architectures for generative sampling of protein structures. Chroma [Ingraham et al. (2022)] trains random graph message passing through roto-translational invariant features, while RFDiffusion [Watson et al. (2022)] fine-tunes pretrained folding model RoseTTAFold [Baek et al. (2021a)] to denoising, employing SE(3)-equivariant transformers in structural decoding [Fuchs et al. (2020)]. [Yim et al. (2023); Anand & Achim (2022)] generalize denoising diffusion to frames of reference, employing Invariant Point Attention [Jumper et al. (2021)] to model three dimensions, while FoldingDiff [Wu et al. (2022a)] explores denoising in angular space. More recently, [Fu et al. (2023)] proposes a latent diffusion model on coarsened representations learned through Equivariant Graph Neural Networks (EGNN) [Satorras et al. (2022)]. In contrast, our model uses roto-translation equivariant features to produce increasingly richer structural representations from autoencoded sequence and coordinates. We propose a novel latent diffusion model that samples directly in this space for generating protein structures.
@@ 3 The Ophiuchus Architecture
We represent a protein as a sequence of $N$ residues each with an anchor position $\mathbf{P}\in\mathbb{R}^{1\times 3}$ and a tensor of irreducible representations of SO(3) $\mathbf{V}^{0:{l_{\max}}}$, where $\mathbf{V}^{l}\in\mathbb{R}^{d\times(2l+1)}$ and degree $l\in[0,{l_{\max}}]$. A residue state is defined as $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})$. These representations are directly produced from sequence labels and all-atom positions, as we describe in the Atom Encoder/Decoder sections. To capture the diverse interactions within a protein, we propose three main components. Self-Interaction learns geometric representations of each residue independently, modeling local interactions of atoms within a single residue. Sequence Convolution simultaneously updates sequential segments of residues, modeling inter-residue interactions between sequence neighbors. Finally, Spatial Convolution employs message-passing of geometric features to model interactions of residues that are nearby in 3D space. We compose these three modules to build an hourglass architecture.
Figure 2: Building Blocks of Ophiuchus: (a) Atom Encoder and (b) Atom Decoder enable the model to directly take and produce atomic coordinates. (c) Self-Interaction updates representations internally, across different vector orders $l$. (d) Spatial Convolution interacts spatial neighbors. (e) Sequence Convolution and (f) Transpose Sequence Convolution communicate sequence neighbors and produce coarser and finer representations, respectively. (g) Hourglass model: we compose those modules to build encoder and decoder models, stacking them into an autoencoder.
@@ 3.1 All-Atom Atom Encoder and Decoder
Given a particular $i$-th residue, let $\mathbf{R}_{i}\in\mathcal{R}$ denote its residue label, $\mathbf{P}_{i}^{\alpha}\in\mathbb{R}^{1\times 3}$ denote the global position of its alpha carbon ($\mathbf{C}_{\alpha}$), and $\mathbf{P}^{\ast}_{i}\in\mathbb{R}^{n\times 3}$ the position of all $n$ other atoms relative to $\mathbf{C}_{\alpha}$. We produce initial residue representations $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})_{i}$ by setting anchors $\mathbf{P}_{i}=\mathbf{P}^{\alpha}_{i}$, scalars $\mathbf{V}_{i}^{l=0}=\textrm{Embed}(\mathbf{R}_{i})$ , and geometric vectors $\mathbf{V}_{i}^{l>0}$ to explicitly encode relative atomic positions $\mathbf{P}^{\ast}_{i}$.
In particular, provided the residue label $\mathbf{R}$, the heavy atoms of most standard protein residues are readily put in a canonical order, enabling direct treatment of atom positions as a stack of signals on SO(3): $\mathbf{V}^{l=1}=\mathbf{P}^{*}$. However, some of the standard residues present two-permutations within their atom configurations, in which pairs of atoms have ordering indices ($v$, $u$) that may be exchanged (Appendix A.1). To handle these cases, we instead use geometric vectors to encode the center $\mathbf{V}^{l=1}_{\textrm{center}}=\frac{1}{2}(\mathbf{P}^{*}_{v}+\mathbf{P}^{*}_{u})$ and the unsigned difference $\mathbf{V}^{l=2}_{\textrm{diff}}=Y_{2}(\frac{1}{2}(\mathbf{P}^{*}_{v}-\mathbf{P}^{*}_{u}))$ between the positions of the pair, where $Y_{2}$ is a spherical harmonics projector of degree $l=2$. This signal is invariant to corresponding atomic two-flips, while still directly carrying information about positioning and angularity. To invert this encoding, we invert a signal of degree $l=2$ into two arbitrarily ordered vectors of degrees $l=1$. Please refer to Appendix A.2 for further details.
@@
Input: All-Atom Relative Positions $\mathbf{P}^{\ast}\in\mathbb{R}^{n\times 3}$
@@
$\mathbf{P}\leftarrow\mathbf{P}^{\alpha}$ $\mathbf{V}^{l=0}\leftarrow\textrm{Embed}(\mathbf{R})$ $\mathbf{V}^{l=1}_{\textrm{ordered}}\leftarrow\textrm{GetOrderablePositions}(\mathbf{R},\mathbf{P}^{\ast})$ ${\mathbf{P}^{*}_{v},\mathbf{P}^{*}_{u}\leftarrow\textrm{GetUnorderablePositionPairs}(\mathbf{R},\mathbf{P}^{\ast})}\;$ $\mathbf{V}^{l=1}_{\textrm{center}}\leftarrow\frac{1}{2}(\mathbf{P}^{*}_{v}+\mathbf{P}^{*}_{u})$ $\mathbf{V}^{l=2}_{\textrm{diff}}\leftarrow Y^{2}\big{(}\frac{1}{2}(\mathbf{P}^{*}_{v}-\mathbf{P}^{*}_{u})\big{)}$ $\mathbf{V}^{l=0:2}\leftarrow\mathbf{V}^{l=1}_{\textrm{ordered}}\oplus\mathbf{V}^{l=1}_{\textrm{center}}\oplus\mathbf{V}^{l=2}_{\textrm{diff}}$ return $(\mathbf{P},\mathbf{V}^{l=0:2})$
@@
Output: All-Atom Relative Positions $\mathbf{P}^{\ast}\in\mathbb{R}^{|\mathcal{R}|\times n\times 3}$
Output: Residue Label Logits $\bm{\ell}\in\mathbb{R}^{|\mathcal{R}|}$
$\mathbf{P}^{\alpha}\leftarrow\mathbf{P}$ $\bm{\ell}\leftarrow{\textrm{LogSoftmax}}\big{(}\textrm{Linear}(\mathbf{V}^{l=0})\big{)}$ $\hat{\mathbf{V}}^{l=1}_{\textrm{ordered}}\leftarrow\textrm{Linear}(\mathbf{V}^{l=1})$ $\hat{\mathbf{V}}^{l=1}_{\textrm{center}},\hat{\mathbf{V}}^{l=2}_{\textrm{diff}}\leftarrow\textrm{Linear}(\mathbf{V}^{l=0:2})$ $\Delta\mathbf{P}_{v,u}\leftarrow\textrm{Eigendecompose}\big{(}\hat{\mathbf{V}}^{l=2}_{\textrm{diff}}\big{)}$ $\mathbf{P}^{*}_{v},\mathbf{P}^{*}_{u}=\hat{\mathbf{V}}^{l=1}_{\textrm{center}}+\Delta\mathbf{P}_{v,u},\hat{\mathbf{V}}^{l=1}_{\textrm{center}}-\Delta\mathbf{P}_{v,u}$ $\mathbf{P}^{\ast}\leftarrow\mathbf{P}^{\ast}\oplus\mathbf{P}^{*}_{v}\oplus\mathbf{P}^{*}_{u}$ return $(\mathbf{P}^{\alpha},\mathbf{P}^{\ast},\bm{\ell})$
@@
This processing makes Ophiuchus strictly blind to ordering flips of permutable atoms, while still enabling it to operate directly on all atoms in an efficient, stacked representation. In Appendix A.1, we illustrate how this approach correctly handles the geometry of side-chain atoms.